OpenStudy (anonymous):

find the height for with the given radius 3cm n volume 350cm3

OpenStudy (earthcitizen):

volume of a cylinder, sphere ?

OpenStudy (anonymous):

cylinder n i must find the height and the surface area.. but they only gave the radius 3cm n volume 350cm3

OpenStudy (earthcitizen):

kk, smartie. How do you think this problem should be solved using the formula given ?

OpenStudy (earthcitizen):

\[V=\pi r ^{2}h\] for cylinder, kk, lv

OpenStudy (earthcitizen):

surface area= \[A=2πr2 + 2πrh = 2πr(r + h)\]

OpenStudy (anonymous):

well wen i try it,it give my problem so can u work it out for me please

OpenStudy (earthcitizen):

yes, I could I would prefer if you understood it so you won't need to worry. Can you make h the subject of the formula in the volume equation ?

OpenStudy (anonymous):

i cnt get the formula out of the volume equation i guess pir2h/2πr2+2πrh=2πr(r+h)

OpenStudy (earthcitizen):

\[h=V/(\pi r ^{2})\]

OpenStudy (anonymous):

tanx so can u work it out for me nw please

OpenStudy (earthcitizen):

lol! Y don't you give it go first. Just plug volume and radius into the equation

OpenStudy (anonymous):

kool loool

OpenStudy (anonymous):

350/pir2

OpenStudy (earthcitizen):

r=3=> r^2 =9

OpenStudy (anonymous):

so datz the height 9cm

OpenStudy (earthcitizen):

h=12.4cm. now you have h and find A, the surface area

OpenStudy (anonymous):

so nw wat is the suface area nw

OpenStudy (earthcitizen):

yh, plug in the values of h=12.4 and r=3 into A=2πr2 + 2πrh = 2πr(r + h) , where ever you h and r replace with values kk

OpenStudy (anonymous):

so it wld be 2pi32 plus 2pi3 12.4=2pir (3 plus 12.4) is datz wat u meant

OpenStudy (earthcitizen):

what is yr pi given as ? 3.14 or 22/7

OpenStudy (anonymous):

3.14

OpenStudy (anonymous):

so nw rewrite the equation using 3.14 as pi for me please

OpenStudy (earthcitizen):

kk, A=2*(3.14)*(3)*(3+12.4)=290.14cm^2

OpenStudy (anonymous):

kool tanx

OpenStudy (earthcitizen):

yh, no p

OpenStudy (anonymous):

kool

OpenStudy (anonymous):

find the height and surface area with the radius of 2cm n volume 350cm3

OpenStudy (earthcitizen):

it's the same steps, just change the numbers

OpenStudy (earthcitizen):

kk, lv!

OpenStudy (anonymous):

can u work it out for me please cuz i am doin another part of the project cuz it is due tomorrow pleasee

OpenStudy (earthcitizen):

\[h=350/((3.14)\times(2)^{2})=350/(3.14\times4)=27.866cm\] \[A=2\pi r(r+h)= 2\times(3.14)\times 2\times(2+27.866)=375.12cm ^{2}\]

OpenStudy (anonymous):

tanx much

OpenStudy (anonymous):

find the height n the surface area with the radius 4cm n volume 350cm3

OpenStudy (anonymous):

please reply